# Optimal Rates for DP-SCO with a Single Epoch and Large Batches

### v1

@@ 4.4 Deferred Proofs
Algorithm 1 with batch size
@@
$B\leq\min\left\{\sqrt{n},\frac{(L+2M\left\|\mathcal{C}\right\|_{2})n^{1.5}\varepsilon}{4\sqrt{2}\left\|\mathcal{C}\right\|_{2}\sqrt{d}\log^{1.5}(T)\sqrt{\ln(4T/\delta)\ln(2.5)/\delta}}\right\}$
@@
and $T=n/B$ (i.e., a single pass over a dataset of size $n$), $\eta_{t}=t+1$, $\beta=M+\frac{(8L+16M\left\|\mathcal{C}\right\|_{2})n^{3/2}}{\left\|\mathcal{C}\right\|_{2}B^{2}}$, $\sigma=\frac{(8\sqrt{2}L+16\sqrt{2}M\left\|\mathcal{C}\right\|_{2})\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ is $(\varepsilon,\delta)$-DP and has the following excess population loss guarantee:
@@
$O\left(\frac{M\left\|\mathcal{C}\right\|_{2}^{2}B^{2}}{n^{2}}+\frac{(L+M\left\|\mathcal{C}\right\|_{2})\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}+\frac{(L+\frac{M\left\|\mathcal{C}\right\|_{2}}{\sqrt{B}})\left\|\mathcal{C}\right\|_{2}\sqrt{d}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right).$
In particular, if we choose $B=\min\{\sqrt{n},Mn^{1.5}\varepsilon/(\sqrt{d}\cdot\text{polylog}(n/B,1/\delta))\}$ then Algorithm 1 requires
$\max\{\sqrt{n},\sqrt{d}\cdot\text{polylog}(n/B,1/\delta)/M\varepsilon\sqrt{n}\}\}$
@@
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left((L\left\|\mathcal{C}\right\|_{2}+M\left\|\mathcal{C}\right\|_{2}^{2})\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right)\right).$
@@
Note that the upper bound on $B$ gives $\beta\geq 16Mn/B\geq 2MT$, satisfying the assumption in Claim 4.2, and also gives $d$ satisfying the assumption in Corollary 4.5. Let $Q=8L+16M\left\|\mathcal{C}\right\|_{2}=O(L+M\left\|\mathcal{C}\right\|_{2})$. We plug the tail bound of Lemma 4.3 into the error bound of Theorem 4.7, use a failure probability of $1/T^{3}$ in that bound (the increase in the expected error due to this low-probability event is $O(L\left\|\mathcal{C}\right\|_{2}/T^{3})$, which is a lower order term in our error bound) and choose $\eta_{t}=t+1,\beta=\frac{QT^{3/2}}{\left\|\mathcal{C}\right\|_{2}\sqrt{B}}$, let:
$c_{DP}=\frac{\sqrt{d\log^{3}(T)\log(T/\delta)\log(1/\delta)}}{\varepsilon}$
@@
We can assume that $d=O(n^{2}\varepsilon^{2})$, i.e. $c_{DP}=O(n\cdot{\rm polylog}\left(T,1/\delta\right)$, as if $d=\omega(n^{2}\varepsilon^{2})$ then any point in $\mathcal{C}$ achieves the desired excess loss bound. Now we split the bound in Theorem 4.7 into six terms and bound them individually. Specifically, from Theorem 4.7 we have:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\underbrace{\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}}_{E_{1}(t)}+\underbrace{\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}}_{E_{2}(t)}+\underbrace{\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}\eta_{0:t}}}_{E_{3}(t)}+\underbrace{\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}}_{E_{4}(t)}+\underbrace{\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}}_{E_{5}(t)}\right)$
$+\underbrace{\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}}_{E_{6}}.$
Now, using (i) $Q/\left\|C\right\|_{2}=O(M)$, (ii) $n=BT$, (iii) $B\leq\sqrt{n}$ (and thus $T\geq\sqrt{n}$) and (iv) $\left\|\mathbf{b}_{t}\right\|_{2}=O\left(\frac{Qc_{DP}}{B\beta}\right)$ by Lemma 4.3 and our choice of $\sigma$, we can bound the contribution of each of $E_{1},E_{2},E_{3},E_{4},E_{5},E_{6}$ individually:
@@
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{1}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right)=O\left(\frac{Q\sqrt{T}(L+\frac{Q}{\sqrt{B}})}{\sqrt{B}\beta}\right)$
$=O\left(\frac{\left\|\mathcal{C}\right\|_{2}(L+\frac{Q}{\sqrt{B}})}{T}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{T}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}\right),$
@@
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{2}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}\right)=O\left(Q\left\|\mathbf{b}_{t}\right\|_{2}\right)=O\left(\frac{Q^{2}c_{DP}}{B\beta}\right)$
$=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{\sqrt{B}T^{3/2}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{T\sqrt{n}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{n}\right),$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{3}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}\eta_{0:t}}\right)=\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}\right),$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{4}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}\right)=O\left(\frac{1}{T^{2}}\sum\limits_{t=0}^{T-1}\frac{Qc_{DP}\left\|\mathcal{C}\right\|_{2}}{B}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{n}\right),$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{5}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)=O\left(\frac{1}{T^{2}}\sum\limits_{t=0}^{T-1}\beta\left(\frac{Qc_{DP}}{B\beta}\right)^{2}\right)=O\left(\frac{Q^{2}c_{DP}^{2}}{B^{2}T\beta}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}^{2}}{B^{3/2}T^{5/2}}\right)$
$=$ $O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}B}{n^{3/2}}{\rm polylog}\left(T,1/\delta\right)\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{n}{\rm polylog}\left(n/B,1/\delta\right)\right),$
$E_{6}=$ $\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}=O\left(\frac{QT^{3/2}}{\left\|\mathcal{C}\right\|_{2}\sqrt{B}}\cdot\frac{1}{T^{2}}\cdot\left\|\mathcal{C}\right\|_{2}^{2}\right)=O\left(\frac{M\left\|\mathcal{C}\right\|_{2}^{2}}{\sqrt{n}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}\right).$
Combining the asymptotic bounds, we get our final bound:
@@
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left(Q\left\|\mathcal{C}\right\|_{2}\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\text{polylog}(n/B,1/\delta)\right)\right).$
@@
4.4.3 Analysis of Stochastic Accelerated Gradient Descent using Independent Gradients

### v2

@@ 5.3 Bounding sensitivity
For $2^{-o(n^{1/6})}\leq\delta\leq 1/n,p\leq\varepsilon^{2}n^{2}$, Algorithm 1 with parameters given by Section 5.2 and using the binary tree mechanism to sample $\mathbf{b}_{i}$ is $(\varepsilon,\delta)$-DP and has the following excess population risk guarantee:
@@
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left(LR\cdot\text{polylog}(n,1/\delta)\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{p}}{\varepsilon n}\right)\right).$
Before we prove the theorem, we note that both the lower bound on $\delta$ and upper bound on $p$ are vacuous: If $p\geq\varepsilon^{2}n^{2}$, then $LR\sqrt{p}/\varepsilon n\geq LR$, and $LR$ is the excess risk achieved by picking any point in $\mathcal{C}$, i.e. the excess risk guarantee in the Theorem is trivially achieved. Similarly, if $\delta=2^{-\Omega(n^{1/6})}$, then $\sqrt{n}\leq{\rm polylog}\left(1/\delta\right)$ and again the excess risk guarantee in the Theorem is trivially achieved.
Proof.
In Lemma 5.5 we chose $\gamma=\delta/4\leq 1/4n$, in which case the high probability events of Theorem 5.1, Lemma 5.4 and Lemma 5.5 happen w.p. at least $1-1/2n$. We give a bound conditioned on these events, which can differ from the unconditional final bound by at most $O(LR/n)$, which is a lower-order term.
From Theorem 5.1, using $Q=2c_{DP}L$, $\left\|\mathbf{b}_{t}\right\|_{2}\leq b_{max}$, and the bound from Lemma 5.4 on $\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}$, simplifying by suppressing terms logarithmic in $n,1/\delta$ we have conditioned on the event of Lemma 5.5:
$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})=$
$\widetilde{O}_{n,1/\delta}\left(\frac{1}{T^{2}}\sum\limits_{t=1}^{T}t^{2}\left(\frac{L\left(\sqrt{M}\left(\sqrt{\frac{LR}{\sqrt{Bt}}}+\sqrt{\frac{\beta b_{max}R}{t}}+\frac{\sqrt{\beta}R}{t}\right)+\frac{L}{\sqrt{Bt}}\right)}{\sqrt{Bt}\beta}\right.\right.$
$\left.\left.+\frac{LR}{\sqrt{B}t^{3/2}}+\frac{\beta b_{max}R}{t^{2}}\right)+\frac{\beta}{T^{2}}R^{2}\right)$
In (21) and (22) we bounded the first term inside the sum, we can reuse the bound to get:
$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})$
$=\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{1}{T^{2}}\sum\limits_{t=1}^{T}t^{2}\left(\frac{LR}{\sqrt{B}t^{1.5}}+\frac{\beta b_{max}R}{t^{2}}\right)+\frac{\beta}{T^{2}}R^{2}\right)$
$=\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{LR}{\sqrt{BT}}+\frac{\beta b_{max}R}{T}+\frac{\beta}{T^{2}}R^{2}\right)$
$=\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{\beta b_{max}R}{T}+\frac{\beta}{T^{2}}R^{2}\right).$
Finally, we plug in the parameter values from (5.2) and use the fact that $\beta b_{max}\leq c_{DP}^{2}L\sqrt{p}/\varepsilon B$ to get:
$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})=\widetilde{O}_{n,1/\delta}\left(LR\left({\sqrt{n}}+\frac{\sqrt{p}}{\varepsilon n}\right)\right).$
5.4 Removing the Assumption $\mathbf{x}^{\dagger}\in\mathcal{C}$
In this section we show how to modify our analysis to the setting where $\mathbf{x}^{\dagger}$ may lie outside the constraint set, i.e. the gradient at $\mathbf{x}^{*}$ is not necessarily 0. In this setting the gradient norm is no longer guaranteed to be decreasing in $t$, which causes our sensitivity to increase and also worsens our final utility bound. To undo both of these effects, we increase $\beta$ to be $\widetilde{\Theta}(MT)$ instead of $\widetilde{\Theta}(M)$, and consequently need to raise $T$ to be $\sqrt{n}$ instead of $\widetilde{\Theta}(n^{1/4})$ to ensure the optimization error $\beta R^{2}/T^{2}$ is $O(1/\sqrt{n})$.
Lemma 5.9 (Bounding $\ell_{2}$-sensitivity).
$\left\|\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)\right\|_{2}\leq 2(Mb_{\max}+2MR+L)\leq 2Mb_{max}+(2c_{M}+1)L.$
Proof.
$\leq\eta_{0:t-1}\left\|\frac{\nabla_{t-1}}{\beta}+\frac{\mathbf{b}_{t-1}}{\eta_{t}}\right\|_{2}+\eta_{t}R$
$\Rightarrow\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{0:t-1}}{\eta_{0:t}}\cdot\frac{\left\|\nabla_{t-1}\right\|_{2}}{\beta}+\frac{\eta_{0:t-1}\left\|\mathbf{b}_{t-1}\right\|_{2}}{\eta_{t}\cdot\eta_{0:t}}+\frac{\eta_{t}}{\eta_{0:t}}R$
$\leq\frac{\sum_{i\leq t-1}\sum_{d\in\mathcal{B}_{i}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta(t+1)B}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}R$
$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}R$
$\Rightarrow\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{t}M\cdot\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+Mb_{\max}+2MR$
$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{\max}+2MR.$ (27)
$\left\|\Delta_{t}(d)\right\|_{2}\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{max}+2MR+L.$ (28)
By induction, this completes the proof of Lemma 5.9. ∎
By using the high-probability upper bound on $\mathbf{b}_{t}$ combined with Lemma 5.9, we can show that $Mb_{max}$ is a lower-order term in Lemma 5.9:
Corollary 5.10.
Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the binary tree mechanism (Lemma 5.6) with $C=\frac{(8c_{M}+4)L}{B\beta}$. Then for $p\leq\frac{B^{2}\beta^{2}\varepsilon^{2}}{32M^{2}\log^{3}(T)\ln(2T/\delta)\ln(2.5/\delta)}$ with probability $21-\delta/$ 2, $\left\|\Delta_{t}(d)\right\|_{2}\leq(8c_{M}+4)L$ for all $d\in supp(\mathcal{D}),t$.
Proof.
The corollary follows from Lemma 5.9 if $b_{max}\leq(2c_{M}+1)L/M$. Plugging $C$ into the bound on $\max_{t}\left\|\mathbf{b}_{t}\right\|_{2}$ from Lemma 5.6, we have:
$b_{max}\leq\frac{(8c_{M}+4)L\log^{3/2}T\sqrt{p\ln(2T/\delta)}}{B\beta\mu}.$
Hence it suffices if:
$\frac{(8c_{M}+4)L\log^{3/2}T\sqrt{p\ln(2T/\delta)}}{B\beta\mu}\leq(2c_{M}+1)L/M.$
Rearranging, we get that it suffices if
$p\leq\left(\frac{(2c_{M}+1)B\beta\mu}{M(8c_{M}+4)\log^{3/2}T\sqrt{\ln(2T/\delta)}}\right)^{2}=\frac{B^{2}\beta^{2}\mu^{2}}{16M^{2}\log^{3}(T)\ln(2T/\delta)}.$
Then by a similar argument to Corollary 5.7, we have:
Corollary 5.11.
Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the binary tree mechanism (Lemma 5.6) with $\sigma=\frac{(16\sqrt{2}c_{M}+8\sqrt{2})L\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ on $\Delta_{i}$. Then for $p\leq\frac{B^{2}\beta^{2}\mu^{2}}{16M^{2}\log^{3}(T)\ln(2T/\delta)}$, Algorithm 1 satisfies $(\varepsilon,\delta)$-DP.
While we require a bound on the dimension to ensure the norm of the noise is bounded, this bound will end up being nearly vacuous: Our choice of $\beta$ is proportional to $n^{3/2}B^{2}$, i.e. the bound we get is roughly $p\leq n^{3}\varepsilon^{2}/B^{2}$. We will choose $B\leq\sqrt{n}$ so it suffices if $p\leq\varepsilon^{2}n^{2}$; this is the range of dimensions where DP-SCO is non-trivial, as for $p>\varepsilon^{2}n^{2}$ any point in $\mathcal{C}$ achieves the (asymptotic) optimal loss bound.
We now give our main result without the assumption $\mathbf{x}^{\dagger}\in\mathcal{C}$:
Theorem 5.12.
@@ 5.4 Removing the Assumption $\mathbf{x}^{\dagger}\in\mathcal{C}$
Algorithm 1 with batch size $B\leq\sqrt{n}$, $T=n/B$, $\eta_{t}=t+1$, $\beta=\frac{(16c_{M}+8)Ln^{3/2}}{RB^{2}}$, $\sigma=\frac{((16\sqrt{2}c_{M}+8\sqrt{2})L)\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ is $(\varepsilon,\delta)$-DP and has the following excess population loss guarantee:
@@
$O\left(\frac{MR^{2}B^{2}}{n^{2}}+\frac{(L+MR)R}{\sqrt{n}}+\frac{(L+\frac{MR}{\sqrt{B}})R\sqrt{p}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right).$
In particular, if we choose $B=\min\{\sqrt{n},Mn^{3/2}\varepsilon/(\sqrt{p}\cdot\text{polylog}(n/B,1/\delta))\}$ then Algorithm 1 requires
$\max\{\sqrt{n},\sqrt{p}\cdot\text{polylog}(n/B,1/\delta)/M\varepsilon\sqrt{n}\}\}$
@@
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left(LR\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{p}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right)\right).$
@@
We assume $p=O\left(\frac{\varepsilon^{2}n^{2}}{{\rm polylog}\left(n,1/\delta\right)}\right)$; if $p=\omega\left(\frac{\varepsilon^{2}n^{2}}{{\rm polylog}\left(n,1/\delta\right)}\right)$, then any point in $\mathcal{C}$ achieves the desired excess risk. Under this assumption, the upper bound on $B$ gives
$\beta\geq 16Mn^{3/2}/B^{2}\geq 16Mn/B\geq 2MT,$
satisfying the assumption in Claim 5.9, and also gives that $d$ satisfies the assumption in Corollary 5.11. Let $Q=O(L)$ in Theorem 5.1 by Corollary 5.10 and $b_{max}=\widetilde{O}_{n,1/\delta}\left(\frac{L\sqrt{p}}{\varepsilon B\beta}\right)$ by Lemma 5.6. We plug the tail bound of Lemma 5.6 into the high-probability error bound of Theorem 5.1, and use a failure probability of $1/T^{2}$ in both Lemma 5.6 and Theorem 5.1 (the increase in the expected error due to this low-probability event is $O(LR/T^{2})$, which is a lower order term in our error bound). Then we have:
$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})$
@@
$=\widetilde{O}_{n,1/\delta}\left(\frac{1}{T^{2}}\sum\limits_{t=1}^{T}t^{2}\left(\frac{L^{2}}{\sqrt{Bt}\beta}+\frac{Lb_{max}}{\sqrt{B}t^{1.5}}+\frac{LR}{\sqrt{B}t^{1.5}}+\frac{\beta b_{max}R}{t^{2}}+\frac{\beta b_{max}^{2}}{t^{2}}\right)+\frac{\beta R^{2}}{T^{2}}\right)$
$=\widetilde{O}_{n,1/\delta}\left(\frac{L^{2}\sqrt{T}}{\sqrt{B}\beta}+\frac{Lb_{max}}{\sqrt{BT}}+\frac{LR}{\sqrt{BT}}+\frac{\beta b_{max}R}{T}+\frac{\beta b_{max}^{2}}{T}+\frac{\beta R^{2}}{T^{2}}\right)$
@@
$=\widetilde{O}_{n,1/\delta}\left(\frac{LRB^{1.5}\sqrt{T}}{n^{3/2}}+\frac{L^{2}\sqrt{p}}{\varepsilon\beta B^{1.5}\sqrt{T}}+\frac{LR}{\sqrt{BT}}+\frac{LR\sqrt{p}}{\varepsilon BT}+\frac{L^{2}p}{\varepsilon^{2}B^{2}\beta T}+\frac{n^{3/2}LR}{B^{2}T^{2}}\right).$
$=\widetilde{O}_{n,1/\delta}\left(\frac{LRB}{n}+\frac{LR\sqrt{Bp}}{\varepsilon n^{3/2}\sqrt{T}}+\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p}}{\varepsilon n}+\frac{LRp}{\varepsilon^{2}n^{3/2}T}+\frac{LR}{\sqrt{n}}\right).$
@@
=$\stackrel{{\scriptstyle(\ast_{1})}}{{=}}\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p}}{\varepsilon n}+\frac{LRp}{\varepsilon^{2}n^{2}}\right).$
=$\stackrel{{\scriptstyle(\ast_{2})}}{{=}}\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p}}{\varepsilon n}\right).$
In $(\ast_{1})$, we use the fact that $B\leq\sqrt{n}$ and $T\geq\sqrt{n}$, and in $(\ast_{2})$ we use the assumption $p=O\left(\frac{\varepsilon^{2}n^{2}}{{\rm polylog}\left(n,1/\delta\right)}\right)$, which implies $p/\varepsilon^{2}n^{2}\leq\sqrt{p}/\varepsilon n$.
@@
5.5 Analysis of Stochastic Accelerated Gradient Descent using Independent Gradients